# Push the Limit of Multi-modal Emotion Recognition by Prompting LLMs with Receptive-Field-Aware Attention Weighting

### v1

@@ Abstract
Introduction
Understanding human emotions is always considered one of the necessary qualities for a powerful AI. Nevertheless, emotion recognition is a tricky task. It requires the models to understand the contents of dialogues, analyze the tones in the acoustic signals, and capture the expressions in the videos. Recent studies (Nguyen et al. 2023; Joshi et al. 2022) have made remarkable progress. These studies focus on how we extract information from each modality and fuse them to better understand the emotion efficiently. However, some dialogues need certain knowledge about modern ethics to infer the correct emotions. One example of this demand is sarcasm, which often has an emotion against the literal meanings. Most current studies only train models from scratch with specific emotion recognition corpora. Although these corpora have thousands of samples, the models are still hard to learn and understand the underlying ethical knowledge.
As shown in Figure 1a, one of the most straightforward strategies to increase the model’s awareness of external knowledge is to integrate a pre-trained natural language model (Wu, Zhang, and Woodland 2021) to pre-process the text modality with its knowledge. These frozen models are pre-trained on large corpora that are not specific for emotion recognition. They could have learned more comprehensive knowledge to understand the dialogues accurately. Previous studies built a vanilla model leveraging the BERT (Devlin 2019) to provide external knowledge for the backbone model for understanding the context. However, the shortage is that given the limited processability and context window of the BERT, only very limited sentences can be sent to BERT simultaneously. Lacking the information of the dialogues could cause the BERT to extract misdirected features.
Large language models have become a new trend in the study of natural language processing. These models with extremely numerous parameters are trained on a great deal of text corpora. They mastered more profound knowledge about the knowledge of our world and society and provided a larger context window compared to the BERT. Regarding the text inputs, hardly any vanilla models can outperform the LLMs. Nonetheless, most of the LLMs, like GPT-3.5, can only process text modality. Even if the latest GPT-4o could process multi-modal inputs, asking LLMs to process these multimedia modalities is costly. For example, a typical video in IEMOCAP (Busso et al. 2008) is 720480, which would cost 425 tokens for one frame when using GPT-4o. A 30-second video with 24 frames per second is over 300k tokens, while the context window size of GPT-4o is only 128k. It is still unacceptable even if we compress it to 8 frames per second with a total of 100k tokens. In addition, from the ablation studies from Nguyen et al. (2023) and Joshi et al. (2022), we find that the improvement brought by the video modality is not significant enough for us to bear the cost. Text is the most important modality in emotion recognition.
@@ Introduction
Lantern poses several challenges. First, we need to deliver information from multimedia modalities to LLMs efficiently. Besides the classification metric used by most of the studies, dimension scores (Russell and Mehrabian 1977) are another option to evaluate emotions. According to the results from Wagner et al. (2023) and Atmaja and Akagi (2020), acoustic signals play a more important role in predicting the dimension scores than linguistic information. Therefore, we deliver emotion classification metrics and dimension scores as supplementary information correlated with multimedia modalities for the LLMs’ reference. Furthermore, multitasking vanilla models are applied for both the categorical metrics and dimensional scores to enable a lightweight pipeline.
@@
Lantern deploys a multi-task vanilla model based on CORECT (Nguyen et al. 2023) and SDT (Ma et al. 2023), while using GPT-4 (Achiam et al. 2023) and Llama-3.1-405B (Dubey et al. 2024) as the LLM agent. A single NVIDIA RTX 3090 GPU is enough to train a vanilla multi-modal model and a lightweight merge model. We conducted experiments on the most commonly used dataset IEMOCAP (Busso et al. 2008). The results confirm that Lantern significantly improves the vanilla model recognition accuracy by 1.23% to 1.80% in 6-way and 4-way settings when using CORECT and 0.79% when using SDT in the 6-way setting.
@@
Related Work
Emotion recognition. The emotion recognition task requires algorithms to analyze the speakers’ emotions from the transcriptions, acoustic signals, and videos. There are two ways to measure the emotion of a speaker: discrete emotions, such as happy, sad, and angry, and emotion dimensions, which describe emotions through arousal, valence, and dominance. IEMOCAP (Busso et al. 2008) provided the transcriptions, acoustic signals, and videos as the input, and both the discrete emotion and dimension labels.
Vanilla multi-modal models. Different modalities can supplement each other by providing more information or confirming shared information to help models better understand the world (Baltrušaitis, Ahuja, and Morency 2018; Majumder et al. 2018). The transformer structure (Vaswani et al. 2017) has significantly improved the performance of fusing multiple modalities. It is often used to handle the visual and language features (Xu, Zhu, and Clifton 2023). Shi and Huang (2023) and Delbrouck et al. (2020) have applied the transformer to this task and achieved great success.
Many researchers choose to use a Graph Neural Network (Gori, Monfardini, and Scarselli (2005), Busso et al. (2008)) to capture the relationship in a conversation. The GNN has different options for its base structure. The most commonly used structures are convolution (Kipf and Welling 2016), recurrence (Nicolicioiu, Duta, and Leordeanu 2019), and attention (Veličković et al. 2018). Nguyen et al. (2023), Joshi et al. (2022), and Hu, Wei, and Huai (2021) deployed GNN to model the conversation to achieve significant results. Hu, Wei, and Huai (2021), Ghosal et al. (2020), Majumder et al. (2019), and Hazarika et al. (2018) use the Gated Recurrent Unit (GRU) as an alternative method.
Though the studies above focused on the classification task, Wagner et al. (2023), Srinivasan, Huang, and Kirchhoff (2022), and Atmaja and Akagi (2020) explored the methods to predict the dimension scores of emotions. Considering the importance of acoustic signals in the task of dimension scores, they applied the wav2vec 2.0 (Baevski et al. 2020) and HuBERT (Hsu et al. 2021) to process the speech.
Prompt LLMs for downstream tasks. The foundation of LLMs is the scaling law (Kaplan et al. 2020), which discovered that the models can achieve remarkable performance improvement by training with an enormously large dataset. For example, Llama 3 (Touvron et al. 2023) is trained on 15T tokens. With a large amount of training data, the LLMs are well-learned to understand the context of conversations with their internal knowledge. Therefore, it is intuitive that many works requiring external knowledge have included LLMs as part of their pipelines. In Visual Question Answering (VQA) tasks, Shao et al. (2023) used vanilla local models to generate heuristics-enhanced prompts to provide information on images for LLMs. Zhang et al. (2023) uses GPT to generate video description and context knowledge for a fine-tuned Llama model for further prediction.
Method
@@
Adaption for Vanilla Models
@@
The classification and dimension scores are both descriptions of the emotion and they are correlative variables (demonstrated in next section). It is trivial to extract the features twice with two different models. Therefore, as shown in Figure 3b, for a given vanilla model, an additional linear layer is added after the backbone to allow it to predict both tasks from the same emotion feature. This multi-task training pattern can reduce the burden at this step and provide a more comprehensive view during the training of the backbone network. The backbone of the vanilla models can be any emotion recognition model. In this work, we adopted the CORECT (Nguyen et al. 2023) as our backbone.
Prompting LLMs
@@
To help the LLMs have a more accurate understanding of the relationship between the dimension scores and the emotion classes, we summarized the statistics of the dimension scores of each emotion (detailed in next section), which is appended in the header for the LLMs’ reference. Inspired by Wei et al. (2022), we selected a series of examples from the training set and manually constructed answers for them. These examples are attached for the LLMs to better understand the description. These answers guide the LLMs to not only examine the meaning of the transcriptions, the values of class probabilities, and the dimension scores but also compare the changes between samples.
@@
Split Strategy
In practice, a dialogue could be too long to fit in the context window of the LLMs. Therefore, we had better split the dialogue into smaller windows. Another reason to split the contents is that the chances of passing the integrity check are negatively related to the number of samples passed to the LLMs. In addition, LLMs have output limits. If we feed too many samples in one prompt, they cannot assign enough thoughts to each sample and reach the limit before finishing adjusting all samples in the prompts.
A simple strategy is to split the dialogue naively and repeat each window times as in Figure 4a. However, a window may require some information from adjacent windows. We can pad dummy samples, like in Figure 4b, so that each window could overlap with each other. Due to the randomness of the prediction results of LLM, we need to fuse the multiple predictions. To build a dimensionally uniform fusion algorithm, each sample will be covered in exactly windows and the paddings will be discarded. Nonetheless, for a sample, its predictions are all based on the same receptive field of a certain length. We hope that the predictions can include as many samples into consideration as possible.
Therefore, we adopted a sliding window method to iterate the dialogue. During sliding, each prediction for the same sample is based on a different receptive field of the dialogue, which achieves feature extraction for the content of dialogues of different lengths. The Figure 4c demonstrates an example when . We can see that the window uses of its maximum length as the step size. It begins with containing only the amount of samples of one step, moves forward one step each time, and ends until the entire window is moved out of the dialogue range. The first and last windows have the least amount of data, we performed padding using dummy samples to prevent their context is not sufficient for LLMs to understand.
@@
Receptive-Field-Aware Weighting
After passing the prompts to the LLMs, we can acquire a new prediction matrix for each sample, where is the number of emotion classes and the represents predictions from the LLMs and one from the vanilla models. When combining these prediction results to make the final decision, the most naive strategy is simply adding up the predicted probabilities (or when ignoring the one produced by the vanilla models). However, this would ignore the internal patterns between each receptive field. We would like to learn the weights for them instead of assigning some arbitrary constants.
Another proposal is to learn an weight matrix and perform an element-wise product before summing up for each class. It ignored the length of the receptive fields for different predictions could be different. The lengths of receptive fields are different for the samples at the beginning or the end of the dialogue. In contrast, the samples in the middle would have the same length for their receptive fields.
Here, we propose a receptive-field-aware attention-driven merge algorithm. is the proportion of the lengths of receptive fields with respect to the dialogue. First, the coarse importance of each prediction is calculated from the lengths of their receptive fields, denoted as ,
@@
Experiments
We evaluated the performance of Lantern on the IEMOCAP (Busso et al. 2008) dataset. Since our method requires the classification label, dimension scores, and transcription, to the best of our knowledge, IEMOCAP is the only dataset that meets these requirements. We first discuss the effectiveness of using dimension scores to supplement information for classification. Then, introduce how we train our model. Finally, the detailed results of micro benchmarks and ablation studies are presented.
Dimension Scores and Emotional Classes
@@
Training Settings
In our experiments, we adopted the CORECT (Nguyen et al. 2023) and SDT (Ma et al. 2023) as the vanilla models. To make these models support multi-tasking, we added 8 lines of code to the network of CORECT, 21 lines to the network of SDT, and around 50 lines to their training procedure to save results for the LLM adjustment. In each modification, we preserved the training method of the classification task. Specifically, in CORECT, a simple linear layer is appended to predict the dimension scores accepting the same feature as the input of the classifier, while, in SDT, self-distillation is also applied to the dimension scores by replacing the KL-diverse with MSE. The gpt-4-0125-preview and Llama-3.1-405B were the LLM deployed in our framework. Our settings completely followed previous studies. The fifth session in IEMOCAP is used as the test set and the validation and training sets are split from the remaining sessions. We train the CORECT model on the training set and select the best model based on the validation set performance. The LLM then makes predictions on both the validation and test sets. The attention-based merge algorithm is trained on the validation set predictions, and the highest-scoring model is used for final test set predictions.
@@
Discussion
As demonstrated in the experiment section, our framework Lantern can effectively improve the performance of the vanilla multi-modal networks. As we demonstrated, any vanilla model, like CORECT (Nguyen et al. 2023) or SDT (Ma et al. 2023), adding an extra prediction head to perform dimension scores and re-trained on the datasets that provided these labels can be deployed in our framework. Similarly, any power LLM can be leveraged as part of Lantern, including both more advanced frozen general LLMs, like GPT and Llama in our experiment, or the LLMs that are specifically fine-tuned for emotion recognition. We believe that the potential of Lantern goes far beyond what has been demonstrated in the experiments. We will keep exploring its potential in the future.

### v2

@@ Abstract
#Corresponding author. E-mails:,
1 Introduction
Understanding human emotions is always considered one of the necessary qualities for a powerful AI. Nevertheless, emotion recognition is a tricky task. It requires the models to understand the contents of dialogues, analyze the tones in the acoustic signals, and capture the expressions in the videos. Recent studies [22, 14] have made remarkable progress. These studies focus on how we extract information from each modality and fuse them to better understand the emotion efficiently. However, some dialogues need certain knowledge about modern ethics to infer the correct emotions. One example of this demand is sarcasm, which often has an emotion against the literal meanings. Most current studies only train models from scratch with specific emotion recognition corpora. Although these corpora have thousands of samples, the models are still hard to learn and understand the underlying ethical knowledge.
Integrate a pre-trained natural language model [33, 17] to pre-process the text modality with its knowledge has been widely used to increase the model’s awareness of external knowledge. As shown in Figure 1a, these frozen models are pre-trained on large corpora that are not specific for emotion recognition. They could have learned more comprehensive knowledge to understand the dialogues accurately. Previous studies built a vanilla model leveraging the BERT [7] to provide external knowledge for the backbone model for understanding the context. However, the shortage is that given the limited processability and context window of the BERT, only very limited sentences can be sent to BERT simultaneously. Lacking the information of the dialogues could cause the BERT to extract misdirected features.
Large language models have become a new trend in the study of natural language processing. These models with extremely numerous parameters are trained on a great deal of text corpora. They mastered more profound knowledge about the knowledge of our world and society and provided a larger context window compared to the BERT. Regarding the text inputs, hardly any vanilla models can outperform the LLMs. Nonetheless, most of the LLMs, like GPT-3.5, can only process text modality. Even if the latest GPT-4o could process multi-modal inputs, asking LLMs to process these multimedia modalities is costly. For example, a typical video in IEMOCAP [5] is 720480, which would cost 425 tokens for one frame when using GPT-4o. A 30-second video with 24 frames per second is over 300k tokens, while the context window size of GPT-4o is only 128k. It is still unacceptable even if we compress it to 8 frames per second with a total of 100k tokens. In addition, from the ablation studies from [22] and [14], we find that the improvement brought by the video modality is not significant enough for us to bear the cost. Text is the most important modality in emotion recognition.
@@ 1 Introduction
Lantern poses several challenges. First, we need to deliver information from multimedia modalities to LLMs efficiently. Besides the classification metric used by most of the studies, dimension scores [24] are another option to evaluate emotions. According to the results from [31] and [2], acoustic signals play a more important role in predicting the dimension scores than linguistic information. Therefore, we deliver emotion classification metrics and dimension scores as supplementary information correlated with multimedia modalities for the LLMs’ reference. Furthermore, multitasking vanilla models are applied for both the categorical metrics and dimensional scores to enable a lightweight pipeline.
@@
Lantern deploys a multi-task vanilla model based on CORECT [22] and SDT [19], while using GPT-4 [1] and Llama-3.1-405B [8] as the LLM agent. A single NVIDIA RTX 3090 GPU is enough to train a vanilla multi-modal model and a lightweight merge model. We conducted experiments on the most commonly used dataset IEMOCAP [5]. The results confirm that Lantern significantly improves the vanilla model recognition accuracy by 1.23% to 1.80% in 6-way and 4-way settings when using CORECT and 0.79% when using SDT in the 6-way setting.
@@
2 Related Work
Emotion recognition. The emotion recognition task requires algorithms to analyze the speakers’ emotions from the transcriptions, acoustic signals, and videos. There are two ways to measure the emotion of a speaker: discrete emotions, such as happy, sad, and angry, and emotion dimensions, which describe emotions through arousal, valence, and dominance. IEMOCAP [5] provided the transcriptions, acoustic signals, and videos as the input, and both the discrete emotion and dimension labels. Vanilla multi-modal models. Different modalities can supplement each other by providing more information or confirming shared information to help models better understand the world [4, 20]. Convolutional Neural Networks (CNNs) [18] has been extensively applied in emotion recognition tasks. The transformer structure [29] has significantly improved the performance of fusing multiple modalities. It is often used to handle the visual and language features [34]. [26] and [6] have applied the transformer to this task and achieved great success.
Many researchers choose to use a Graph Neural Network ([10], [5]) to capture the relationship in a conversation. The GNN has different options for its base structure. The most commonly used structures are convolution [16], recurrence [23], and attention [30]. [22], [14], and [13] deployed GNN to model the conversation to achieve significant results. [13], [9], [21], and [11] use the Gated Recurrent Unit (GRU) as an alternative method.
Though the studies above focused on the classification task, [31], [27], and [2] explored the methods to predict the dimension scores of emotions. Considering the importance of acoustic signals in the task of dimension scores, they applied the wav2vec 2.0 [3] and HuBERT [12] to process the speech. Prompt LLMs for downstream tasks. The foundation of LLMs is the scaling law [15], which discovered that the models can achieve remarkable performance improvement by training with an enormously large dataset. For example, Llama 3 [28] is trained on 15T tokens. With a large amount of training data, the LLMs are well-learned to understand the context of conversations with their internal knowledge. Therefore, it is intuitive that many works requiring external knowledge have included LLMs as part of their pipelines. In Visual Question Answering (VQA) tasks, [25] used vanilla local models to generate heuristics-enhanced prompts to provide information on images for LLMs. [35] uses GPT to generate video description and context knowledge for a fine-tuned Llama model for further prediction.
3 Method
@@
3.1 Adaption for Vanilla Models
@@
The classification and dimension scores are both descriptions of the emotion and they are correlative variables (demonstrated in next section). It is trivial to extract the features twice with two different models. Therefore, as shown in Figure 3b, for a given vanilla model, an additional linear layer is added after the backbone to allow it to predict both tasks from the same emotion feature. This multi-task training pattern can reduce the burden at this step and provide a more comprehensive view during the training of the backbone network. The backbone of the vanilla models can be any emotion recognition model. In this work, we adopted the CORECT [22] as our backbone.
3.2 Prompting LLMs
@@
To help the LLMs have a more accurate understanding of the relationship between the dimension scores and the emotion classes, we summarized the statistics of the dimension scores of each emotion (detailed in next section), which is appended in the header for the LLMs’ reference. Inspired by [32], we selected a series of examples from the training set and manually constructed answers for them. These examples are attached for the LLMs to better understand the description. These answers guide the LLMs to not only examine the meaning of the transcriptions, the values of class probabilities, and the dimension scores but also compare the changes between samples.
@@
3.3 Split Strategy
In practice, a dialogue could be too long to fit in the context window of the LLMs. Therefore, we had better split the dialogue into smaller windows. Another reason to split the contents is that the chances of passing the integrity check are negatively related to the number of samples passed to the LLMs. In addition, LLMs have output limits. If we feed too many samples in one prompt, they cannot assign enough thoughts to each sample and reach the limit before finishing adjusting all samples in the prompts. A simple strategy is to split the dialogue naively and repeat each window times as in Figure 4a. However, a window may require some information from adjacent windows. We can pad dummy samples, like in Figure 4b, so that each window could overlap with each other. Due to the randomness of the prediction results of LLM, we need to fuse the multiple predictions. To build a dimensionally uniform fusion algorithm, each sample will be covered in exactly windows and the paddings will be discarded. Nonetheless, for a sample, its predictions are all based on the same receptive field of a certain length. We hope that the predictions can include as many samples into consideration as possible. Therefore, we adopted a sliding window method to iterate the dialogue. During sliding, each prediction for the same sample is based on a different receptive field of the dialogue, which achieves feature extraction for the content of dialogues of different lengths. The Figure 4c demonstrates an example when . We can see that the window uses of its maximum length as the step size. It begins with containing only the amount of samples of one step, moves forward one step each time, and ends until the entire window is moved out of the dialogue range. The first and last windows have the least amount of data, we performed padding using dummy samples to prevent their context is not sufficient for LLMs to understand.
@@ 3.3 Split Strategy
3.4 Receptive-Field-Aware Weighting
After passing the prompts to the LLMs, we can acquire a new prediction matrix for each sample, where is the number of emotion classes and the represents predictions from the LLMs and one from the vanilla models. When combining these prediction results to make the final decision, the most naive strategy is simply adding up the predicted probabilities (or when ignoring the one produced by the vanilla models). However, this would ignore the internal patterns between each receptive field. We would like to learn the weights for them instead of assigning some arbitrary constants. Another proposal is to learn an weight matrix and perform an element-wise product before summing up for each class. It ignored the length of the receptive fields for different predictions could be different. The lengths of receptive fields are different for the samples at the beginning or the end of the dialogue. In contrast, the samples in the middle would have the same length for their receptive fields. Here, we propose a receptive-field-aware attention-driven merge algorithm. is the proportion of the lengths of receptive fields with respect to the dialogue. First, the coarse importance of each prediction is calculated from the lengths of their receptive fields, denoted as ,
@@ 3.4 Receptive-Field-Aware Weighting
4 Experiments
We evaluated the performance of Lantern on the IEMOCAP [5] dataset. Since our method requires the classification label, dimension scores, and transcription, to the best of our knowledge, IEMOCAP is the only dataset that meets these requirements. We first discuss the effectiveness of using dimension scores to supplement information for classification. Then, introduce how we train our model. Finally, the detailed results of micro benchmarks and ablation studies are presented.
4.1 Dimension Scores and Emotional Classes
@@
4.2 Training Settings
In our experiments, we adopted the CORECT [22] and SDT [19] as the vanilla models. To make these models support multi-tasking, we added 8 lines of code to the network of CORECT, 21 lines to the network of SDT, and around 50 lines to their training procedure to save results for the LLM adjustment. In each modification, we preserved the training method of the classification task. Specifically, in CORECT, a simple linear layer is appended to predict the dimension scores accepting the same feature as the input of the classifier, while, in SDT, self-distillation is also applied to the dimension scores by replacing the KL-diverse with MSE. The gpt-4-0125-preview and Llama-3.1-405B were the LLM deployed in our framework. Our settings completely followed previous studies. The fifth session in IEMOCAP is used as the test set and the validation and training sets are split from the remaining sessions. We train the CORECT model on the training set and select the best model based on the validation set performance. The LLM then makes predictions on both the validation and test sets. The attention-based merge algorithm is trained on the validation set predictions, and the highest-scoring model is used for final test set predictions.
@@ 4.4 Ablation Studies
5 Discussion
As demonstrated in the experiment section, our framework Lantern can effectively improve the performance of the vanilla multi-modal networks. As we demonstrated, any vanilla model, like CORECT [22] or SDT [19], adding an extra prediction head to perform dimension scores and re-trained on the datasets that provided these labels can be deployed in our framework. Similarly, any power LLM can be leveraged as part of Lantern, including both more advanced frozen general LLMs, like GPT and Llama in our experiment, or the LLMs that are specifically fine-tuned for emotion recognition. We believe that the potential of Lantern goes far beyond what has been demonstrated in the experiments. We will keep exploring its potential in the future.
@@ 6 Conclusion
Acknowledgment
This work is supported in part by National Natural Science Foundation of China (No.61936015), Natural Science Foundation of Shanghai (No.24ZR1430600), XJTLU Research Development Funding (No.RDF-21-02-072) and Shanghai Key Laboratory of Trusted Data Circulation and Governance, and Web3.